OpenStudy (anonymous):

omg i will die .. please somebody help me .. i have a home work about lines and line segments , i don not understant , i am in eleven grade i need somebody want video call .

OpenStudy (rational):

ask ur question here

OpenStudy (anonymous):

I have alot of questions, and theres lines and line segments

OpenStudy (rational):

start by posting ur first question

OpenStudy (anonymous):

line means :a long, narrow mark or band. "a row of closely spaced dots will look like a continuous line" synonyms: dash, rule, bar, score; More line segments means : In geometry, a line segment is a part of a line that is bounded by two distinct end points, and contains every point on the line between its end points. Examples of line segments include the sides of a triangle or square. ...

OpenStudy (anonymous):

thank you very much , its a great help. can you please help me with the first question ???

OpenStudy (anonymous):

dats alot of them

OpenStudy (anonymous):

just the first question :(

OpenStudy (rational):

finding distance on number line is easy : just subtract the coordinates

OpenStudy (rational):

look at the given number line, can you tell me the coordinates of A and C ? A = ? C = ?

OpenStudy (rational):

btw the doggie in ur dp is cute :3

OpenStudy (anonymous):

-9 ?? thank you very much that is incredible and this's my dog its so cute thank you

OpenStudy (rational):

you're almost correct keep this in mind : distances or lengths can never be NEGATIVE.

OpenStudy (rational):

A is located at -8 C is located at -1 is that right ? the pic is bit blurred... can u confirm :)

OpenStudy (anonymous):

yes

OpenStudy (rational):

A is located at -8 C is located at +1 right ?

OpenStudy (anonymous):

yep

OpenStudy (rational):

C is +1 or -1 ? which one is it ?

OpenStudy (anonymous):

oh +1

OpenStudy (rational):

Ok so we have : A = -8 C = 1 so the length AC = C - A = 1- -8 = 1 + 8 = 9

OpenStudy (rational):

if u get negative numbers, just ignore the sign.

OpenStudy (rational):

what about the length BD ?

OpenStudy (anonymous):

9

OpenStudy (rational):

Very good, its same as AC's length so we say AC is congruent to BD

OpenStudy (anonymous):

thank you :) :)

OpenStudy (anonymous):

i am so happy

OpenStudy (anonymous):

if the distance isn't the same, is it not congruent?

OpenStudy (rational):

np you're good in geometry i see :) see if u can do 2nd question

OpenStudy (anonymous):

you're smart in geometry too what grade are you in?

OpenStudy (rational):

geometry is indeed fun :) once u know about lines and angles u wil find it easy to deal with ! im done with studies wbu ? :)

OpenStudy (anonymous):

i am so tired i have done other homeworks could you help me with this geometry homework tomorrow please if u have time? im in 11th grade. what hour are you going to be connected tomorrow? :( :( :( :( :(

OpenStudy (rational):

il be online mostly, drop me a msg when u get on have good sleep :) gnite !

OpenStudy (anonymous):

oh my god thank you very much youre a good person youre a great person youre the first person who wants to help me because nobody wants to help me with this.............................. i am going to sleep right now sweet dreams for you